OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1

OpenStudy (anonymous):

A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5

OpenStudy (anonymous):

I believe it's D because -4 and -1 is -5

OpenStudy (amistre64):

if (x-a) is a factor of P(x) we would expect: \[P(x)=(x-a)Q(x)\] \[P(a)=(\underbrace{a-a}_{=0})Q(a)\]

OpenStudy (amistre64):

we can divide out the roots ...

OpenStudy (campbell_st):

well if x = -4 then x + 4 is a factor and x =-1 means x + 1 is a factor so the quadratic is f(x) = (x + 4)(x + 1) now just expand it

OpenStudy (amistre64):

x4 + 5x3 + 5x2 + 5x - 5 0 -4 -4 -4 -4 -4 1 1 1 1 -9 ^^ not 0, -4 aint a root of D

OpenStudy (amistre64):

x4 + 5x3 + 5x2 + 5x + 4 0 -4 -4 -4 -4 -4 1 1 1 1 0 1 1 1 1 0 -1 0 -1 -1 1 0 1 0 x^2+1 is whats left after we factor out the other 2 roots.

OpenStudy (anonymous):

@amistre64 can you explain what you mean with a lot less numbers please?

OpenStudy (amistre64):

prolly not, its take more words than numbers. follow campbells approach

OpenStudy (amistre64):

D doesnt have -4 as a root, so D is out

OpenStudy (amistre64):

test it out for yourself

OpenStudy (anonymous):

@amistre64 =(x+4)(x+1) =(x)(x)+(x)(1)+(4)(x)+(4)(1) =x^2+x+4x+4 =x^2+5x+4???

OpenStudy (anonymous):

@amistre64 does that look right?

OpenStudy (amistre64):

no, xx is a root of 0

OpenStudy (amistre64):

wait, i didnt read it correctly

OpenStudy (amistre64):

(x+4)(x+1) = x^2 + 5x + 4 what we want is 2 more facotrs that are not real ... some (x-(a + bi))(x-(a-bi))

OpenStudy (amistre64):

multiplying it all out is going to e a bother, and prolly not that useful. the division of the real roots i think is more beneficial

OpenStudy (amistre64):

now if we look ahead a bit, we would want some other factor to be x^2 + n for some n>0 so that it has no real roots left in it (x^2 + 5x + 4)(x^2 + n) x^2(x^2 + n) + 5x(x^2 + n) + 4(x^2 + n) x^4 +n x^2 + 5x^3 + 5n x + 4x^2 + 4n x^4 + 5x^3 +n x^2+ 4x^2 +5n x + 4n x^4 + 5x^3 +(4+n) x^2+ 5n x + 4n now we can compare this part by part to see what options fit

OpenStudy (anonymous):

Start at the last one you wrote and evaluate? @amistre64

OpenStudy (amistre64):

yeah, its not D, so the last term has to be either +4 or -4 let n=+1 or -1 and see what matches

OpenStudy (amistre64):

lol, we said n cant be less than 0 .. soo n=1 has to be it

OpenStudy (anonymous):

Okay. So n =1 I got x4+5x3+(4+1)(x2)+(5)(1)x+(4)(1) x4+5x3+(4+1)(x2)+(5)(1)x+(4)(1) =x4+5x3+5x2+5x+4 @amistre64

OpenStudy (amistre64):

thats what we are looking for

OpenStudy (anonymous):

@amistre64 I can't believe I figured it out. Thank you so much for all your help. You are a savior

OpenStudy (amistre64):

good luck :)